Account Change in Inventories Private Consumption Spending Interest Payments made by private firms Government Subsidies Net Private Investment (excludes depreciation) 45 Government Purchases of Goods Depreciation Wages & Salaries (pre-tax) Government Purchases of Services Government Transfer Payments Imports Exports Sales Tax Divedend Payments & Retained Earnings ValuePart of 2000 365 30 200 つつ 1100 90 400 400 500 95 900 (a) Calculate the value of GDP using the expenditure approach (b) Calculate the value of GDP using the income approach. (c) If the table above contains all the relevant income and expenditure accounts for an economy in a given year, explain why your answers to (a) and (b) should be equal d) Calculate the value of non-factor payments
Answer #1

Answer:

a) GDP at Expenditure Method = Consumption + Investment + Governemnt Expenditure + Net Trade

Consumption = 2000

Investment = 45

Governemnt Expenditure = Governemnt Purchase of Goods + Governemnt Purchase of Services = 200+90 = 290

Net Trade = Export - Import = 500 - 400 = 100

Change in inventories = -5

Hence GDP = 2000+45+290 -5+100 = 2430

b) GDP at Income method = Wages & Salary + Interest Income + Rent Earning + Tax - Subsidy

  Wages & Salary = 1100

Interest Income = 365

Dividend Payment & Retained Earnings = 900

Sales Tax = 95

Governemnt Subisidy = 30

Hence GDP = 1100+365+900+95-30 = 2430

c) GDP gives an account of all final goods and services that are produced in an economy for a given year. The income and expenditure method are two ways of measuring the final goods and services that are produced. Expenditure method looks into how much is spend while generating the same goods and services and Income approach take care of how much income is given to different economic agents for producing the same number of goods and services for a given time period.

d) Value of Non-Factor Payments = Value of Governemnt Transfer Payments = 400